OpenStudy (anonymous):

Help??? On a certain hot summer day, 482 people used the public swimming pool. The daily prices are $1.25 for children and $2.25 for adults. The receipts for admission totaled $865.25. How many children and how many adults swam at the public pool that day?

OpenStudy (chaise):

Let C be the number of children entering the pool Let A be the number of adults entering the pool A+C=482 1.25C+2.25A=865.25 Solve simultaneously.

OpenStudy (anonymous):

what do you mean by solve simultaneously?

OpenStudy (mathstudent55):

It means to solve as a system of equations.

OpenStudy (anonymous):

can you help me do that? Q~Q

OpenStudy (chaise):

A+C=482 Eqn(1) 1.25C+2.25A=865.25 Eqn(2) Rearrange Eqn(1) to give: C=482-A Eqn(3) Substitute Eqn(3) in to Eqn(2) 1.25(482-A)+2.25A=865.25 Expand, collect like terms and solve for A, resub back in to Eqn(1) to find C.

OpenStudy (anonymous):

OMG you are so difficult to understand like what the heck is Eqn supposed to stand for equation ? and i ONLY thought there were two parts of solving this type of problem what the heck MRS.DORTCH IS GONNA HAVE MY A** -.-